# Landscape of Correlated Orders in Strained Bilayer Nickelate Thin Films

### v2

@@ Abstract
Then we investigate the instabilities in the NiO2 bilayer slightly away from the interface, which is influenced solely by epitaxial strain. The orbital-resolved band structure is depicted in Fig.3(a), showing two pockets in the absence of charge doping. The disappearance of the $d_{z^{2}}$-bonding $\gamma$ hole pocket is attributed to the reduced onsite energy of $d_{z^{2}}$ orbital due to the compressive strain. However, this hole pocket can emerge through a Lifshitz transition when the hole doping level $\delta$ exceeds $\delta_{c}=0.22$. Fig. 3(b) illustrates the Fermi surfaces in such a case with $\delta=0.3$. Prior to the Lifshitz transition, our FRG calculations show that the system predominantly exhibits SDW instability with either no doping or slight hole doping according to the phase diagram in Fig. 3(d). However, after the Lifshitz transition (black dashed line), superconductivity instability becomes robust and can even persist under strong Hund’s coupling, as shown in Fig. 3(d). A representative FRG flow for the case with a hole doping $\delta=0.3$ is displayed in Fig. 3(c), where $s_{\pm}$-superconductivity is dominant, with the gap function depicted in the inset. This pairing state is primarily attributed to the interlayer pairing, which induces a sign change in superconducting gaps between the bonding and anti-bonding states. This behavior is analogous to that observed in the bulk under pressure Sakakibara et al. (2024a); Gu et al. (2025); Yang et al. (2023a); Liu et al. (2023). These suggest that the emergence of the $\gamma$ hole pocket around the M point has a significant effect on the SDW order and $s_{\pm}$-wave superconductivity, which is related to Fermi surface nesting. Without the $\gamma$ pocket, the nesting between $\alpha$ and $\beta$ pockets contributes to significant SDW fluctuation at $\bm{q}_{4}$, inducing a dominant SDW instability. However, the appearance of the $\gamma$ pocket introduces an additional strong nesting between $\beta$ and $\gamma$ pockets at $\bm{q}_{2}$, competing with SDW fluctuation at $\bm{q}_{4}$. This is apparent from the FRG flow in the SDW channel shown in Fig. 3(c) and this competition suppresses SDW instability. Remarkably, both Fermi surface nestings occur between bonding and anti-bonding pockets and thus generate strong interlayer antiferromagnetic fluctuations (see SM), significantly promoting interlayer pairing and stabilizing $s_{\pm}$-wave superconductivity. Consequently, the observed superconductivity in compressed thin films is likely to occur in NiO2 bilayers slightly away from the interface, where the substrate can contribute to both strain effect and hole doping due to the substitution of La/Pr with Sr Zhou et al. (2025).
@@
Enhanced SDW instability in stretched LNO thin films. We further investigate the instability of LNO thin films under tensile strain. Experimental results indicate that LNO films on SrTiO3 do not exhibit superconductivity but rather show insulating behavior Bhatt et al. (2025), despite theoretical calculations suggesting that their Fermi surfaces are similar to those of the bulk under pressure Geisler et al. (2025); Zhao and Botana (2025). Fig. 4(a) presents the ab initio band structures in a tetragonal phase with an in-plane lattice constant of $a=3.90$ , consistent with previous calculations Geisler et al. (2025); Zhao and Botana (2025). The corresponding Fermi surfaces from an effective TB model are shown in Fig. 4(b), where all three pockets become more rectangular. It is evident that the $\alpha$ pocket incorporates more $d_{z^{2}}$ orbital, while the contribution of the $d_{x^{2}-y^{2}}$ orbital to the $\beta$ pocket around the M point is also enhanced. These generate enhanced Fermi surface nesting at $\bm{q}_{2}$ and $\bm{q}_{4}$ vectors. More intriguingly, the rectangular $\alpha$ pocket aligns well with the $\gamma$ pocket when shifted by $\bm{q}_{5}=(\pi,\pi)$, indicating strong Fermi surface nesting between them. The bare susceptibility along the high-symmetry path are displayed in Fig. 4(c) for $\delta=0$ and $\delta=\pm 0.1$. A sharp nesting peak around $\bm{q}_{5}$ is dominant, while the peaks at $\bm{q}_{4}$ and $\bm{q}_{2}$ are also prominent. Electron doping can relatively enhance nesting at $\bm{q}_{2,4}$. Due to the competition among multiple types of spin fluctuations, the undoped system does not develop instability for a low $J_{H}/U<0.1$ above the cutoff of $10^{-4}$, as shown in Fig. 4(d). However, when Hund’s coupling is increased, the system exhibits a robust SDW instability with a wavevector of $\bm{q}_{4}$, regardless of hole or electron doping, as shown in Fig. 4(d). The nesting at $\bm{q}_{5}$ involves two bonding pockets and induces interlayer ferromagnetic coupling, which suppresses interlayer spin-singlet pairing. This, together with significant SDW fluctuations at $\bm{q}_{2}$ and $\bm{q}_{4}$, renders superconductivity mostly subdominant. The absence of superconductivity and robust SDW seem to be consistent with the experimental observation of insulating behavior in the resistivity of stretched LNO thin films Bhatt et al. (2025).
[…]
Discussion and conclusion. Our calculations highlight the pivotal role of the hole pocket around the M point in promoting superconductivity in bilayer nickelates under compressive strain, as depicted in the schematic phase diagram in Fig. 5(a). This is intimately related to both its high density of states and bonding state nature. Its appearance introduces an additional Fermi surface nesting with vectors $\bm{q}_{2}$, and leads to frustration of SDW fluctuations, thus destabilizing the SDW instability. Meanwhile, two Fermi surface nestings involving vectors $\bm{q}_{2}$ and $\bm{q}_{4}$ occur between bonding and anti-bonding pockets, generating significant interlayer antiferromagnetic coupling, as shown in Fig. 5(a) and (b). When SDW instabilities are not dominant, their fluctuations will promote interlayer pairing, leading to the emergence of superconductivity. This scenario is consistent with the absence of superconductivity for the NiO2 layer at the interface and under tensile strain. In the former case, electron doping can lead to a Lifshitz transition, where an antibonding electron pocket appears around the $\Gamma$ point. Its nesting with the antibonding $\beta$ pocket contributes to ferromagnetic interlayer coupling, thus suppressing interlayer spin-singlet pairing. In the latter case, the strong nesting at $\bm{q}_{5}$ between two bonding pockets also contributes to ferromagnetic interlayer coupling, hindering interlayer pairing.
The maximum $T_{c}$ observed in strained thin films is approximately 50 K Liu et al. (2025); Zhou et al. (2025), which is significantly lower than that of the bulk under pressure. This value could potentially be increased with future improvements in sample quality. However, our calculations suggest that the $T_{c}$ in strained thin films may consistently remain lower than in the bulk. This is due to the elongation of the out-of-plane Ni-Ni distance under compressive strain and Sr doping, which weakens interlayer antiferromagnetic coupling and interlayer pairing. Therefore, further external pressure can enhance interlayer coupling and raise the transition temperature of strained thin films, similar to the infinite-layer nickelates Wang et al. (2022). Under tensile strain, strong Fermi surface nesting leads to competing SDW fluctuations and the system develops an SDW instability. The Fermi surface nesting can be suppressed by charge doping, potentially allowing superconductivity to emerge when the strong SDW order is diminished.
In summary, we investigate the electronic structures and competing correlated instabilities of strained LNO thin films on substrates to identify the key factors for achieving superconductivity. We find that the compressed NiO2 bilayer at the interface, which is heavily electron-doped from the substrate, is unlikely to exhibit superconductivity. In contrast, the NiO2 bilayer slightly away from the interface displays density-wave instability when undoped or slightly hole-doped. However, when this bilayer is moderately hole-doped, resulting in the emergence of a hole pocket around the M point, two SDW orders compete, destabilizing the SDW instability. Intriguingly, both SDWs exhibit interlayer antiferromagnetic coupling, and their fluctuations promote interlayer pairing, leading to robust $s_{\pm}$-wave superconductivity. This may account for the superconductivity observed in thin films. For the stretched NiO2 bilayer, a robust SDW instability is observed over a wide range of interaction parameters due to significantly enhanced Fermi surface nesting, despite the presence of a hole pocket around the M point. We further discuss the differences between bulk and thin films and strategies to enhance $T_{c}$. Our work establishes a unified scenario for superconductivity in both pressurized bulk and strained thin films of bilayer nickelates.
@@
Acknowledgments. We acknowledge the supports by the Ministry of Science and Technology (Grant No. 2022YFA1403901), National Natural Science Foundation of China (No. 11920101005, No. 11888101, No. 12047503, No. 12322405, No. 12104450) and the New Cornerstone Investigator Program. X.W. is supported by the National Key R&D Program of China (Grant No. 2023YFA1407300) and the National Natural Science Foundation of China (Grants No. 12447103). C.C.L. is supported by the RIKEN TRIP initiative (RIKEN Quantum).
Note added. After completing this work, we became aware of an independent study on superconductivity in thin films Shao et al. (2025). This study focuses on the pairing symmetry in the Amam phase without the $\gamma$ hole pocket around the M point, using an either-band tight-binding model. In contrast, our work systematically explores both superconductivity and charge density wave states for compressed and stretched thin films across various levels of charge doping, highlighting the crucial role of the hole pocket around the M point in promoting superconductivity.

### v3

@@ Abstract
Next, we investigate the instabilities in the NiO2 bilayer slightly away from the interface, which is influenced solely by epitaxial strain and Sr-diffusion-induced doping Zhou et al. (2025). The orbital-resolved band structure is depicted in Fig.3(a), showing two pockets in the absence of hole doping. The $d_{z^{2}}$ bonding band is located below $E_{\text{F}}$ due to its reduced onsite energy. However, with hole doping, this band approaches $E_{\text{F}}$, and eventually crosses $E_{\text{F}}$, resulting a Lifshitz transition when $\delta$ exceeds $\delta_{c}=0.22$. Fig. 3(b) illustrates the Fermi surfaces in such a case with $\delta=0.3$. Our FRG calculations show that, in the undoped and slight-hole-doped cases, the system exhibits prominent SDW instability for strong Hund’s coupling, while $s_{\pm}$ superconductivity occurs at weak Hund’s coupling. However, when the $d_{z^{2}}$-bonding band approaches $E_{\text{F}}$ (incipient) or crosses $E_{\text{F}}$ to generate a Lifshitz transition (black dashed line), superconductivity instability becomes enhanced and robust, and can even persist under strong Hund’s coupling, as shown in Fig. 3(d). After the Lifshitz transition, Tc drops monotonically with increasing doping. A representative FRG flow for the case with a hole doping $\delta=0.3$ is displayed in Fig. 3(c), where $s_{\pm}$-superconductivity is dominant, with the gap function depicted in the inset. This pairing state is primarily attributed to the interlayer pairing on $d_{z^{2}}$ orbital. These results suggest that when the $d_{z^{2}}$ bonding band becomes incipient or crosses $E_{F}$ to form the $\gamma$ pocket, it significantly suppresses SDW order and generates robust $s_{\pm}$-wave superconductivity, which is related to competing Fermi surface nesting.
This is intimately related to the interlayer magnetic coupling, which can be characterized by the bare interlayer susceptibility, defined as,
$\chi^{0}_{t\alpha,b\beta}(\bm{q},i\omega_{n})=-\frac{1}{N}\sum_{\bm{k}\nu\nu^{\prime}}a^{b\beta}_{\nu}(\bm{k})a^{t\alpha*}_{\nu}(\bm{k})a^{t\alpha}_{\nu^{\prime}}(\bm{k}+\bm{q})$
$\times a^{b\beta*}_{\nu^{\prime}}(\bm{k}+\bm{q})\frac{n_{F}[\epsilon_{\nu}(\bm{k})]-n_{F}[\epsilon_{\nu^{\prime}}(\bm{k}+\bm{q})]}{i\omega_{n}+\epsilon_{\nu}(\bm{k})-\epsilon_{\nu^{\prime}}(\bm{k}+\bm{q})}.$ (3)
Here $\nu/\nu^{\prime}$ is the band index, $n_{F}(\epsilon)$ is the Fermi distribution function, $a^{l\alpha}_{\nu}(\bm{k})$ is the $l\alpha$-th component of the eigenvector for band $\nu$ by diagonalizing the TB model and $\epsilon_{\nu}(\bm{k})$ is the corresponding eigenvalue. Nesting between two bands with the same mirror parity, namely both bonding or antibonding, contributes the even-channel susceptibility $\chi^{0,e}$ Bötzel et al. (2024). This yields a positive product of eigenvector elements in the above formula, resulting a positive $\chi^{0}_{t\alpha,b\beta}$ and, consequently, ferromagnetic interlayer coupling. In contrast, nesting between a bonding and an antibonding band contributes the odd-channel susceptibility $\chi^{0,o}$, and yields a negative product and a negative $\chi^{0}_{t\alpha,b\beta}$, leading to AFM interlayer coupling. When the $d_{z^{2}}$ band resides away from $E_{F}$, the nesting between bonding $\alpha$ and antibonding $\beta$ pockets contributes to significant SDW fluctuation at $\bm{q}_{4}$, inducing a dominant SDW instability. However, when the $d_{z^{2}}$-bonding flat band becomes incipient Chen et al. (2015); Mishra et al. (2016); Nakata et al. (2017) or crosses $E_{F}$ to generate the $\gamma$ pocket around the M point, this introduces an additional strong SDW fluctuation with the nesting vector $\bm{q}_{2}$ connecting $\beta$ and $\gamma$ pockets, competing with SDW fluctuation at $\bm{q}_{4}$. This is apparent from the FRG flow in the SDW channel shown in Fig. 3(c), where the close competition between two SDWs with different wave vectors suppresses the SDW instability. Remarkably, both Fermi surface nestings occur between bonding and anti-bonding bands and thus generate strong interlayer AFM fluctuations, collectively promoting interlayer pairing and stabilizing $s_{\pm}$-wave superconductivity. The presence of $\gamma$ pocket introduces strong interlayer FM coupling from $\gamma-\gamma$ nesting  Ryee et al. (2024), suppressing interlayer pairing and resulting a Tc reduction with increasing doping. Consequently, the observed superconductivity in compressed thin films is likely to occur in NiO2 bilayers slightly away from the interface, where the substrate can contribute to both strain effect and Sr-diffusion-induced hole doping Zhou et al. (2025).
@@
Enhanced SDW instability in stretched LNO thin films. We further investigate the instability of LNO thin films under tensile strain. Experimental results indicate that LNO films on SrTiO3 do not exhibit superconductivity but rather show insulating behavior Bhatt et al. (2025), despite theoretical calculations suggesting that their Fermi surfaces are similar to those of the bulk under pressure Geisler et al. (2025); Zhao and Botana (2025). Under tensile strain, LNO exhibit orthorhombic distortion Bhatt et al. (2025), but its electronic structure remains similar to the tetragonal phase (see SM). Here, we present the band structure (Fig. 4(a)) for the tetragonal phase with using an in-plane lattice constant of $a=3.90$ Å , consistent with previous results Geisler et al. (2025); Zhao and Botana (2025). The corresponding Fermi surfaces from an effective TB model (Fig. 4(b)) show all three pockets becoming more rectangular. The $\alpha$ pocket gains increased $d_{z^{2}}$ character, and the $d_{x^{2}-y^{2}}$ contribution to the $\beta$ pocket is also enhanced, leading to strong Fermi surface nesting at $\bm{q}_{2}$ and $\bm{q}_{4}$. Intriguingly, the rectangular $\alpha$ pocket aligns well with the $\gamma$ pocket when shifted by $\bm{q}_{5}=(\pi,\pi)$, indicating strong Fermi surface nesting between them. From the susceptibility at $\delta=0$ shown in Fig. 4(c), we observe a dominant sharp nesting peak around $\bm{q}_{5}$ in the even channel, with weaker peaks at $\bm{q}_{4}$ and $\bm{q}_{2}$ in the odd channel. Electron doping can suppress the nesting at $\bm{q}_{5}$. Due to multiple competing spin fluctuations, the undoped system does not develop instability for a low $J_{H}/U<0.1$ above the cutoff of $10^{-4}$, as shown in Fig. 4(d). However, when Hund’s coupling is increased, the system exhibits a robust SDW instability with a wavevector of $\bm{q}_{4}$, regardless of hole or electron doping. The nesting at $\bm{q}_{5}$ involves two bonding pockets and induces interlayer ferromagnetic coupling, which suppresses interlayer pairing. This, together with $\beta-\beta$ and $\gamma-\gamma$ nesting, renders superconductivity mostly subdominant. The absence of superconductivity and robust SDW seem to be consistent with the experimental observation of insulating behavior in the resistivity of stretched LNO thin films Bhatt et al. (2025).
Discussion and conclusion. Our results reveal a unique superconducting mechanism in bilayer nickelates, governed by scatterings between Fermi surfaces with opposite mirror parity. Unlike iron based superconductors, where different SDW fluctuations generate competing pairings Hirschfeld et al. (2011), in LNO two SDW fluctuations cooperatively enhance interlayer $s_{\pm}$-wave pairing through shared interlayer AFM coupling. In contrast, SDW fluctuations from the same-parity nesting are pairing-breaking. When pressure or charge doping causes the $d_{z^{2}}$ bonding band to become incipient or cross $E_{\text{F}}$, competing magnetic fluctuations destabilize SDW instability and superconductivity emerges. This mechanism offers a unified scenario for superconductivity in both bulk and thin-film LNO. Once the bonding $\gamma$ pocket fully appears, it introduces both pair-forming and pair-breaking effects. Our calculations show that this suppression dominates beyond the Lifshitz transition, producing a $T_{c}$ maximum near the transition (Fig. 1), which is consistent with the superconducting dome in Sr-doped LNO thin films Hao et al. (2025).
The maximum $T_{c}$ reported in strained thin films is about 50 K Liu et al. (2025); Zhou et al. (2025), well below the bulk under pressure. While improved sample quality may raise this value, our results suggest $T_{c}$ in strained films will remain lower because compressive strain weakens the interlayer AFM coupling and thus interlayer pairing. External pressure along c axis should counteract this effect, strengthening interlayer coupling and increasing $T_{c}$, analogous to infinite-layer nickelates Wang et al. (2022). Under tensile strain, strong pair-breaking nesting suppresses superconductivity; however, charge doping can weaken this effect and may enable superconductivity once the SDW is suppressed.
In summary, we establish a direct link between ordering tendencies and Fermiology in bilayer nickelates: scattering between Fermi surfaces of different mirror parities governs SDW and superconductivity. Robust $s_{\pm}$-wave superconductivity can emerge when the $d_{z^{2}}$ bonding band becomes incipient or crosses $E_{\text{F}}$. These results provide a unified framework for superconductivity in both pressurized bulk and strained thin films.
@@
Acknowledgments. We acknowledge the supports by National Natural Science Foundation of China (No. 12494594, No.11920101005, No. 11888101, No. 12047503, No. 12322405, No. 12104450), the Ministry of Science and Technology (Grant No. 2022YFA1403901), and the New Cornerstone Investigator Program. X.W. is supported by the National Key R&D Program of China (Grant No. 2023YFA1407300) and the National Natural Science Foundation of China (Grants No. No. 12574151, 12447103 and 12447101). C.C.L. is supported by the RIKEN TRIP initiative (RIKEN Quantum).